OpenStudy (anonymous):

A 5-mile cab ride costs $7.20. A 9-mile cab ride costs $11.60. Find a linear equation that models a relationship between cost c and distance d. c = 1.44d + 4.40 d = 1.10c + 4.40 c = 1.29d + 1.70 c = 1.10d + 1.70

OpenStudy (anonymous):

let cost be a function of distance so (cost,distance) we have two points : (5,7.20) , (9,11.60) using two points.. can you calculate slope ?

OpenStudy (anonymous):

slope = 1.1 ?

OpenStudy (anonymous):

very good! lets call the slope m. now use one of the points and get the line equation since we treat y as c (cost) and x as d (distance) the line equation is given by c = c1 + m(d-d1) where d1,c1 are the coordinates of one of your points

OpenStudy (anonymous):

So, would it be written like this? c = 7.20 + 1.1(d-9)

OpenStudy (anonymous):

you have to choose one of the points : (5,7.20) or (9,11.60) you dont mix between them

OpenStudy (anonymous):

I don't know which ones to choose?

OpenStudy (anonymous):

it doesnt matter .. pick one point and it will be your (d1,c1)

OpenStudy (anonymous):

eventually you will get the same answer.. no matter which of the point you choose

OpenStudy (anonymous):

So , it will be C = 7.20 + 1.1(d-5)

OpenStudy (anonymous):

yes ! its good

OpenStudy (fellowroot):